OpenStudy (anonymous):

the half-life f a radioactive element is 149 day, but your sample will not be useful to you after 70% of the radioactive nuclei orignally present have disintegrated. about how many days can you use the sample? a. 224 bb. 259 c. 264 d. 269

OpenStudy (anonymous):

pretty much identical to this one http://openstudy.com/users/danielle2012#/updates/4f71b5bfe4b07738f5ade82a

OpenStudy (anonymous):

\[(\frac{1}{2})^{\frac{t}{149}}=.3\] \[\frac{t}{149}=\frac{\ln(.3)}{\ln(.5)}\] \[t=\frac{149\ln(.3)}{\ln(.5)}\]